OpenStudy (314dokg):

SOLID MENSURATION QUESTION (Need help ASAP) I've been trying different kinds of method but it's not even helping A piece of wire length 52m is cut into two parts. Each part is then bent to form a square. It is found that the combined area of the two squares is 109 sq. m. Find the measures of the sides of the two squares. Need help ASAP :( tA = ((52 - Ps1)/4 )^2 + ((52 - Ps2)/4)^2 i tried this but it didnt work

OpenStudy (mrnood):

if you cut the wire into two lengths x & y then you know that x+y = 52 the squares have sides x/4 and y/4 so the area of the squares are (x/4)^2 and (y/4)^2 so (x/4)^2 +(y/4)^2 = 109 I have solved by graphing (plot both equations and find intersection points), but not yet solved analytically

OpenStudy (314dokg):

@MrNood Soo how can i solve the sides? the x and y's... it seems impossible if thats the given equation.

OpenStudy (mrnood):

you have 2 equations x+y =52 (x/4)^2 +(y/4)^2 = 109 from the first y= (52-x) substitiute that into the second equation you will get a quadratic in x solve for x then solve for y in the first equation. The answer is x/4 and y/4 for th esquare side lenghts

OpenStudy (314dokg):

Oh i see i get it now

OpenStudy (314dokg):

let me try it first

OpenStudy (mrnood):

note - the quadratic has TWO real solutions you can choose EITHER of them since we don't care whch 'end' of the wire is which

OpenStudy (314dokg):

Thank you so much @MrNood XD

OpenStudy (mrnood):

np. and oyur answer is?....

OpenStudy (314dokg):

first square is 10m per sides. with the perimeter of 40m second square is 3m per sides with the perimeter of 12m then total perimeter is 40 + 12 = 52 and area is 10^2 + 3^2 = 100 + 9 = 109m^2 i finally get it now thanks a lot

OpenStudy (mrnood):

well done good to see someone stick it out through some complicated calculation!

OpenStudy (314dokg):

took me 1 week to understand the question properly... i used this site as my last resort. glad i finally understand this. lol thanks again